# Jarník-type theorem for self-similar sets

### v1

@@ Abstract.
Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular measure $\mu$. For any $\tau>1/d$ and any ‘inhomogeneous’ vector ${\bm{\theta}}\in\mathbb{R}^{d}$, let $W_{d}(\psi_{\tau},{\bm{\theta}})$ denote the set of $(\psi_{\tau},{\bm{\theta}})$-well approximable numbers, where $\psi_{\tau}(q)=q^{-\tau}$. Assuming a local estimate for the $\mu$-measure of the intersections of $K$ with the neighborhoods of ‘rational’ vectors $(\mathbf{p}+{\bm{\theta}})/q$, we establish a sharp upper bound for the Hausdorff dimension of $K\cap W_{d}(\psi_{\tau},{\bm{\theta}})$, together with some nontrivial lower bounds when $\tau$ is below a certain threshold. One of the lower bounds becomes sharp in the one-dimensional homogeneous case ($d=1,\theta=0$) for a class of sufficiently thick self-similar sets $K$, and moreover $K\cap W_{1}(\psi_{\tau},0)$ has full $(\delta+\frac{2}{1+\tau}-1)$-Hausdorff measure. These results have several applications:
@@ 1. Introduction
Although Mahler himself did not pose a more specific question, this nonetheless inspired subsequent studies on the metric properties of the intersection of $W_{d}(\psi,{\bm{\theta}})$ with self-similar sets, of which the Cantor set is a special case. Recent years have witnessed significant progress in this direction, particularly with the establishment of Khintchine-type theorems for self-similar measure and Besicovitch–Jarník-type theorem for missing digits sets. In the following, we discuss the problem in two parts, focusing separately on the Lebesgue measure and the Hausdorff dimension.
For the Lebesgue measure, Kleinbock, Lindenstrauss and Weiss [22] established that if $\mu$ is a friendly measure—which includes strongly irreducible self-similar measures—then $\mu(W_{d}(\psi_{\tau},\mathbf{0}))=0$ for all $\tau>1/d$. These results were later extended by Pollington and Velani [28] to arbitrary functions $\psi$ for which a certain $\mu$-volume sum converges. Building on this foundation, Einsiedler, Fishman, and Shapira [14], as well as Simmons and Weiss [30], demonstrated that $\mu(W_{d}(\varepsilon\psi_{1/d},\mathbf{0}))$ has full measure for any $\varepsilon>0$, with $\mu$ being either a self-similar measure or an ergodic measure. A complete Khintchine-type theorem for a large class of self-similar measures was established by Khalil and Luethi [20], as well as by Datta and Jana [13], using different techniques. The case of all strongly irreducible self-similar measures was finally settled by Bénard, He and Zhang [3, 2]:
@@
For the Hausdorff dimension, comparatively little is known, and the corresponding problems appear to be at least as challenging as those encountered in the context of Khintchine-type theorem for self-similar measures. Levesley, Salp and Velani [26] first established that, for any missing digits set $K$, the Hausdorff dimension of its intersection with $W_{1}(\psi_{\tau},0)$ is at least $(\operatorname{\dim_{H}}K)/(1+\tau)$. In fact, the Cantor-type subset constructed in their proof only guarantees that points are well-approximated by rational points lying within the missing digits set $K$ itself, which are far fewer than rational points in $\mathbb{R}$. Therefore, this lower bound is far from optimal. They further posed the following conjecture, which has recently been fully resolved by Chen [10]. Define the set of very well approximable points by
@@
Let $\mu_{1/3}$ be the restriction of $\mathcal{H}^{\frac{\log 2}{\log 3}}$ to $K_{1/3}$. It is currently unclear whether $\dim_{l^{1}}\mu_{1/3}>1/2$, and consequently (1.1) in Theorem Y cannot be applied to $K_{1/3}$. Although the result for the Hausdorff dimension of $K_{1/3}\cap\operatorname{VWA}_{1}(0)$ has already been established by Chen [10], Yu’s approach still has its merits, and with suitable refinements, it can be extended to the inhomogeneous setting (see Theorem 1.5 (1)). The second conclusion (1.2) in Theorem Y does not hold for $K_{1/3}$, since $\dim_{l^{1}}\mu_{1/3}\leq\operatorname{\dim_{H}}K_{1/3}$ (see [31, Lemma 1.4 (1) and §3.6]), and hence $\operatorname{\dim_{H}}K_{1/3}\dim_{l^{1}}\mu_{1/3}\leq(\frac{\log 2}{\log 3})^{2}\approx 0.63093^{2}<1/2$. Later, Chow, Varjú and Yu [11] showed that (1.1) and (1.2) in Theorem Y can be respectively applied to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
@@
The following theorem provides a sharp upper bound and a reasonably good lower bound for the Hausdorff dimension of well-approximable points on a compact set $K\subset\mathbb{R}^{d}$ whose measure, supported on $K$, satisfies $(\alpha,\beta,{\bm{\theta}})$-local counting property.
@@
Theorem 1.4 (1) extends Theorem C to higher dimensions. Moreover, for sufficiently thick self-similar sets, we are able to obtain the Hausdorff measure of the set $K\cap E_{1}(\tau,0)$, which generalizes Theorem Y on missing digits sets to thick self-similar sets. However, Theorem 1.4 (2) cannot be straightforwardly extended to higher dimensions, as explained in Remarks 2 and 8.
@@
Let $K\subset\mathbb{R}$ be a missing digits set, and let $\mu$ be the restriction of $\mathcal{H}^{\operatorname{\dim_{H}}K}$ to $K$. Then, the following statements hold.
@@
Item (1) of Theorem 1.5 provides an affirmative answer to a question posed by Yu [31, Remark 9.4] concerning inhomogeneous Diophantine approximation on missing digits sets. Item (2) goes further by strengthening the conclusion of Theorem Y from Hausdorff dimension to Hausdorff measure. As mentioned earlier, it follows from the work of Chow, Varjú, and Yu [11, Propositions 2.4 and 2.5] that items (1) and (2) apply respectively to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
@@
In Section 2, we introduce a useful tool for proving Theorem 1.2. Sections 3 and 4 are devoted to proving Theorem 1.2, while Section 5 presents the proof of Corollary 1.3. In Section 6, we focus on two specific classes of sets: self-similar sets and missing digits sets, which are treated in two separate subsections. More precisely, we verify that these two classes satisfy the $(\alpha,\beta,{\bm{\theta}})$-local counting property for some parameters $\alpha$, $\beta$ and ${\bm{\theta}}\in\mathbb{R}^{d}$, respectively, which in turn yields Theorems 1.4 and 1.5.
@@ 2. Hausdorff measure and content
Proposition 2.6 (Mass distribution principle [6, Lemma 1.2.8]).
@@
(4.6) $\mu\big(B(\mathbf{x},r)\cap A_{Q}(cQ^{-\tau},{\bm{\theta}})\big)\asymp\mu\big(B(\mathbf{x},r)\big)\cdot Q^{1-\tau d}\asymp r^{\delta}\cdot Q^{1-\tau d}.$
@@
Case 2: $Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}\leq r<Q^{-\beta}$. Then, applying the estimate from Case 1 with $r=Q^{-\beta}$, we obtain
@@ 4.1. Estimate of the $\nu$-measure of balls for general cases
In [23, Proposition 6.3], Kleinbock and Weiss proved that any $\delta$-Ahlfors regular measure with $\delta>d-1$ is absolutely $(C,\delta+1-d)$-decaying for some constant $C$ depending only on $d$ and the measure $\mu$, in the following sense: for any ball $B(\mathbf{x},r)\subset\mathbb{R}^{d}$, any hyperplane $L\subset\mathbb{R}^{d}$, and any $\varepsilon>0$,
@@
Case 3B: $Q^{-(1+1/d)}/(8d!)\leq r<Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$. Divide the ball $B(\mathbf{x},r)$ into smaller hypercubes of equal side length $Q^{-(1+1/d)}/(8d!)$. If any such hypercube $\mathcal{C}$ has a nonempty intersection with $K$, then $2\mathcal{C}\cap K$ contains a ball of radius $|\mathcal{C}|$, where $2\mathcal{C}$ denotes the hypercube with the same center as $\mathcal{C}$ but twice the side length. Since any point in $K$ belongs to at most $3^{d}$ different sets of the form $2\mathcal{C}\cap K$, a simple volume argument shows that the ball $B(\mathbf{x},r)$ can intersect at most
@@
hypercubes $\mathcal{C}$. Furthermore, since each $\mathcal{C}$ is convex and its $d$-dimensional Lebesgue measure is bounded by $Q^{-(1+d)}/(8d!)^{d}\leq(2Q)^{-(1+d)}/(d!)$, it follows from the simplex lemma above (see Lemma 4.1) that the rational points with denominator $1\leq q\leq 2Q$ contained in $\mathcal{C}$ lie in some hyperplane $L\subset\mathbb{R}^{d}$. Therefore,
(4.8) $\mathcal{C}\cap A_{Q}(cQ^{-\tau},\mathbf{0})\subset\mathcal{C}\cap L^{(cQ^{-(1+\tau)})}.$
@@
Case 4B: $cQ^{-(1+\tau)}\leq r<Q^{-(1+1/d)}/(8d!)$. Since the $d$-dimensional Lebesgue measure of the ball $B(\mathbf{x},r)$ is less than $Q^{-(1+d)}/(4d!)^{d}$, by the same reason as (4.8),
@@
Case 5B: $0<r<cQ^{-(1+\tau)}$. Then,
@@
$\frac{1-\delta^{2}-\beta\delta}{(1-\delta)^{2}}\leq\frac{1-\delta^{2}-(2-2\delta)\delta}{(1-\delta)^{2}}=1,$
@@ 6. Some Examples
Let $\Lambda^{*}=\bigcup_{k\geq 1}\Lambda^{k}$ denote the set of finite words over $\Lambda$. Let $k\geq 1$. Given a word ${\bm{\omega}}=(\omega_{1},\dots,\omega_{k})\in\Lambda^{k}$, we define
(6.3) $\begin{split}&\phi_{\bm{\omega}}:=\phi_{\omega_{1}}\circ\cdots\circ\phi_{\omega_{k}},\quad\rho_{\bm{\omega}}:=\rho_{{\bm{\omega}}_{1}}\cdots\rho_{{\bm{\omega}}_{k}},\\ &K_{\bm{\omega}}:=\phi_{\bm{\omega}}(K)\quad\text{and}\quad\mu_{\bm{\omega}}:=\frac{\mu|_{K_{\bm{\omega}}}}{\mu(K_{\bm{\omega}})}=\mu\circ\phi_{\bm{\omega}}^{-1}.\end{split}$
We call $\mu_{\bm{\omega}}$ a branch of $\mu$. Intuitively, $\mu_{\bm{\omega}}$ represents the measure $\mu$ restricted to a ‘cylinder’ set $K_{\bm{\omega}}$ and then rescaled to a probability measure.
@@
For $t>0,\mathbf{x}\in\mathbb{R}^{d}$, let
(6.4) $g(t)=\begin{pmatrix}e^{t/d}\mathbb{I}_{d}&0\\ 0&e^{-t}\end{pmatrix}\text{ and }u(\mathbf{x})=\begin{pmatrix}\mathbb{I}_{d}&\mathbf{x}\\ 0&1\end{pmatrix}$,
@@
Corollary 6.2 ([10, Corollary 2.6]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $f\in B_{p,l}^{\infty}(X)$. Then, for any ${\bm{\omega}}\in\Lambda^{*}$ and $y\in X$, we have
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|\ll\mathrm{inj}(y)^{-1}\rho_{\bm{\omega}}^{-1-\frac{d\kappa}{d+1}}\mathcal{S}_{p,l}(f)e^{-\kappa t},$
@@
Since $\mu_{\bm{\omega}}=\mu\circ\phi_{\bm{\omega}}^{-1}$ and $\phi_{\bm{\omega}}(\mathbf{x})=\rho_{\bm{\omega}}\mathbf{x}+\mathbf{b}_{\bm{\omega}}$ with $\mathbf{b}_{\bm{\omega}}=\phi_{\bm{\omega}}(\bf 0)$, we have
$\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})$
@@
$=$ $\int_{\mathbb{R}^{d}}f\left(g\left(t+\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{x})g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right){\,\mathrm{d}}\mu(\mathbf{x}),$
@@
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\bm{\omega}}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|$
@@
$\mathrm{inj}(hz)\geq C_{1}\,\|\mathrm{Ad}(h^{-1})\|_{\infty}\,\mathrm{inj}(z),$
where $\mathrm{Ad}$ denotes the adjoint action of $G$ on $\mathfrak{g}$ and $\|\cdot\|_{\infty}$ is the operator norm. Applying this with $h=g(-\frac{d}{d+1}\ln\rho_{\bm{\omega}})u(\mathbf{b}_{\bm{\omega}})$ yields
$\begin{split}\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right)&\geq C_{1}\left\|\operatorname{Ad}\left(g\left(\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)\right)\right\|_{\infty}\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y)\\ &\geq C_{1}\cdot\rho_{\bm{\omega}}^{-1}\,\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y).\end{split}$
Since $u(\mathbf{b}_{\bm{\omega}})$ is uniformly bounded over ${\bm{\omega}}\in\Lambda^{*}$, there exists a constant $C_{2}>0$ such that $\mathrm{inj}(u(\mathbf{b}_{\bm{\omega}})y)\geq C_{2}\,\mathrm{inj}(y)$. Combining these bounds, we obtain
$\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\bm{\omega}}\right)u(\mathbf{b}_{\bm{\omega}})y\right)^{-1}\leq C_{1}^{-1}C_{2}^{-1}\rho_{\bm{\omega}}^{-1}\mathrm{inj}(y)^{-1},$
@@
Lemma 6.3 ([20, Theorem 9.1 and Lemma 12.7]).

### v2

@@ Abstract.
Let $K\subset\mathbb{R}^{d}$ be a compact subset equipped with a $\delta$-Ahlfors regular measure $\mu$. For any $\tau>1/d$ and any “inhomogeneous” vector ${\bm{\theta}}\in\mathbb{R}^{d}$, let $W_{d}(\psi_{\tau},{\bm{\theta}})$ denote the set of $(\psi_{\tau},{\bm{\theta}})$-well approximable numbers, where $\psi_{\tau}(q)=q^{-\tau}$. Assuming a local estimate for the $\mu$-measure of the intersections of $K$ with the neighborhoods of “rational” vectors $(\mathbf{p}+{\bm{\theta}})/q$, we establish a sharp upper bound for the Hausdorff dimension of $K\cap W_{d}(\psi_{\tau},{\bm{\theta}})$, together with some nontrivial lower bounds when $\tau$ is below a certain threshold. One of the lower bounds becomes sharp in the one-dimensional homogeneous case ($d=1,\theta=0$) for a class of sufficiently thick self-similar sets $K$, and moreover $K\cap W_{1}(\psi_{\tau},0)$ has full $(\delta+\frac{2}{1+\tau}-1)$-Hausdorff measure. These results have several applications:
@@
We also provide some non-trivial missing digits set $K\subset[0,1]^{d}$ whose intersection with $W_{d}(\psi_{\tau},\mathbf{0})$ has full $(\delta+\frac{1+d}{1+\tau}-d)$-Hausdorff measure.
@@ 1. Introduction
Although Mahler himself did not pose a more specific question, this nonetheless inspired subsequent studies on the metric properties of the intersection of $W_{d}(\psi,{\bm{\theta}})$ with self-similar sets, of which the Cantor set is a special case. Recent years have witnessed significant progress in this direction, particularly with the establishment of Khintchine-type theorems for self-similar measures and Besicovitch–Jarník-type theorems for missing digits sets. In the following, we discuss the problem in two parts, focusing separately on the self-similar measure and the Hausdorff dimension.
For the self-similar measure, Kleinbock, Lindenstrauss and Weiss [27] established that if $\mu$ is a friendly measure—which includes strongly irreducible self-similar measures—then $\mu(W_{d}(\psi_{\tau},\mathbf{0}))=0$ for all $\tau>1/d$. These results were later extended by Pollington and Velani [37] to arbitrary functions $\psi$ for which a certain $\mu$-volume sum converges. Building on this foundation, Einsiedler, Fishman and Shapira [17], as well as Simmons and Weiss [39], demonstrated that $\mu(W_{d}(\varepsilon\psi_{1/d},\mathbf{0}))$ has full measure for any $\varepsilon>0$, with $\mu$ being either a self-similar measure or an ergodic measure. A complete Khintchine-type theorem for a large class of self-similar measures was established by Khalil and Luethi [25], as well as by Datta and Jana [15], using different techniques. The case of all strongly irreducible self-similar measures was finally settled by Bénard, He and Zhang [3, 2]:
@@
For the Hausdorff dimension, comparatively little is known, and the corresponding problems appear to be at least as challenging as those encountered in the context of Khintchine-type theorems for self-similar measures. Levesley, Salp and Velani [32] first established that, for any missing digits set $K$, the Hausdorff dimension of its intersection with $W_{1}(\psi_{\tau},0)$ is at least $(\operatorname{\dim_{H}}K)/(1+\tau)$. In fact, the Cantor-type subset constructed in their proof only guarantees that points are well-approximated by rational points lying within the missing digits set $K$ itself. This type of problem is also referred to as intrinsic Diophantine approximation, which focuses on approximating points in a fractal set using rational points contained within the set rather than arbitrary rational points in $\mathbb{Q}$. For more related research, see, for example [16, 19, 33, 35, 40].
Since the rational points contained within the missing digits set $K$ are far fewer than those in $\mathbb{R}$, the lower bound $(\operatorname{\dim_{H}}K)/(1+\tau)$ is far from optimal. Levesley, Salp and Velani [32] further posed the following conjecture, which has recently been fully resolved by Chen [12]. Define the set of very well approximable points by
@@
Let $\mu_{1/3}$ be the restriction of $\mathcal{H}^{\frac{\log 2}{\log 3}}$ to $K_{1/3}$. It is currently unclear whether $\dim_{l^{1}}\mu_{1/3}>1/2$, and consequently (1.1) in Theorem Y cannot be applied to $K_{1/3}$. Although the result for the Hausdorff dimension of $K_{1/3}\cap\operatorname{VWA}_{1}(0)$ has already been established by Chen [12], Yu’s approach still has its merits, and with suitable refinements, it can be extended to the inhomogeneous setting (see Theorem 1.5 (1)). The second conclusion (1.2) in Theorem Y does not hold for $K_{1/3}$, since $\dim_{l^{1}}\mu_{1/3}\leq\operatorname{\dim_{H}}K_{1/3}$ (see [42, Lemma 1.4 (1) and §3.6]), and hence $\operatorname{\dim_{H}}K_{1/3}\dim_{l^{1}}\mu_{1/3}\leq(\frac{\log 2}{\log 3})^{2}\approx 0.63093^{2}<1/2$. Later, Chow, Varjú and Yu [13] showed that (1.1) and (1.2) in Theorem Y can be respectively applied to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
@@
The following theorem provides a sharp upper bound and a reasonably good lower bound for the Hausdorff dimension of well-approximable points on a compact set $K\subset\mathbb{R}^{d}$ whose measure, supported on $K$, satisfies the $(\alpha,\beta,{\bm{\theta}})$-local counting property.
@@
Theorem 1.4 (1) extends Theorem C to higher dimensions. Moreover, for sufficiently thick self-similar sets, we are able to obtain the Hausdorff measure of the set $K\cap E_{1}(\tau,0)$, which generalizes Theorem Y on missing digits sets to thick self-similar sets. However, Theorem 1.4 (2) cannot be straightforwardly extended to higher dimensions, as explained in Remarks 2 and 9.
@@
Let $K\subset[0,1]$ be a missing digits set, and let $\mu$ be the restriction of $\mathcal{H}^{\operatorname{\dim_{H}}K}$ to $K$. Then, the following statements hold.
@@
Theorem 1.5 (1) provides an affirmative answer to a question posed by Yu [42, Remark 9.4] concerning inhomogeneous Diophantine approximation on missing digits sets. Theorem 1.5 (2) goes further by strengthening the conclusion of Theorem Y from Hausdorff dimension to Hausdorff measure. As mentioned earlier, it follows from the work of Chow, Varjú, and Yu [13, Propositions 2.4 and 2.5] that Theorem 1.5 (1) and (2) apply respectively to 5-ary and 7-ary missing digits sets, each with exactly one missing digit.
As noted in Remarks 2 and 4, the $(\alpha,\beta,\mathbf{0})$-local counting property alone is insufficient to determine the Hausdorff dimension of $K\cap W_{d}(\psi_{\tau},\mathbf{0})$ for $d>1$. It is therefore natural to ask whether there exists a non-trivial fractal set $K$ for which the Hausdorff dimension of $K\cap W_{d}(\psi_{\tau},\mathbf{0})$ is exactly $\delta+\frac{1+d}{1+\tau}-d$. The following result shows that certain non-trivial missing digits sets provide a positive answer to this question.
Theorem 1.6.
Let $K\subset[0,1]^{d}$ be a missing digits set satisfying the following assumptions:
  1. (A1)
    $K=K_{1}\times\cdots\times K_{d-1}\times[0,1]$, where $K_{j}$ ($1\leq j\leq d-1$) is a missing digits set;
  2. (A2)
    there exists $\frac{d}{d+1}<\gamma<1$ such that, for each $1\leq j\leq d-1$, the restriction of $\mathcal{H}^{\operatorname{\dim_{H}}K_{j}}$ to $K_{j}$, denoted by $\mu_{j}$, satisfies $\dim_{l^{1}}\mu_{j}>\gamma$.
Then, there exists $\alpha>1/d$ such that
$\mathcal{H}^{\delta+\frac{1+d}{1+\tau}-d}\big(K\cap E_{d}(\tau,\mathbf{0})\big)=\infty,\quad\text{for any $\tau\in(1/d,\alpha)$.}$
Remark 8.
The key ingredient of Theorem 1.6 can be described as follows. It is important to note that the classical mass transference principle (MTP for short) of Beresnevich and Velani [5] cannot be applied directly in this setting. The main obstacle is that some rational vectors $\mathbf{p}/q$ may lie outside $K$. In particular, it can happen that $B(\mathbf{p}/q,r)\cap K\neq\varnothing$ while $B(\mathbf{p}/q,r^{\tau})\cap K=\varnothing$ for any $\tau>1$ unless $K=[0,1]^{d}$. A genuine way to bypass this difficulty is to shrink the ball not in every direction, but only in, for example, the $d$-th direction. This motivates the requirement that the $d$-th component of $K$ be the full interval $[0,1]$, so that shrinking a ball in this direction maintains a substantial intersection with $K$. More precisely, let $K$ be a missing digits set satisfying (A1), and let $\prod_{j=1}^{d}B(\frac{p_{j}}{q},r_{j})\subset[0,1]^{d}$. If $K\cap\prod_{j=1}^{d}B(\frac{p_{j}}{q},r_{j})\neq\varnothing$, then
$K\cap\left(\Bigg(\prod_{j=1}^{d-1}B\bigg(\frac{p_{j}}{q},r_{j}\bigg)\Bigg)\times B\bigg(\frac{p_{d}}{q},r_{d}^{\tau}\bigg)\right)\neq\varnothing\quad\text{for any $\tau>1$},$
since the $d$-th component of $K$ is $[0,1]$. This strategy allows us to adapt the idea of the MTP without the stronger assumption that $K=[0,1]^{d}$. However, because the original MTP applies only to $\limsup$ sets defined by balls, the approach we adopt here is closer in spirit to the MTP from rectangles to small open sets [21, Theorem 2.11], as established by the first author.
@@
In Section 2, we introduce a useful tool for proving Theorems 1.2 and 1.6. Sections 3 and 4 are devoted to proving Theorem 1.2, while Section 5 presents the proof of Corollary 1.3. In Section 6, we focus on two specific classes of sets: self-similar sets and missing digits sets, which are treated in two separate subsections. More precisely, we verify that these two classes satisfy the $(\alpha,\beta,{\bm{\theta}})$-local counting property for some parameters $\alpha$, $\beta$ and ${\bm{\theta}}\in\mathbb{R}^{d}$, respectively, which in turn yields Theorems 1.4 and 1.5. In Section 7, we prove Theorem 1.6.
@@ 2. Hausdorff measure and content
Proposition 2.6 (Mass distribution principle [8, Lemma 1.2.8]).
@@
[…]
@@
Case 2: $Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}\leq r<Q^{-\beta}$. Applying the estimate from Case 1 with $r=Q^{-\beta}$, we obtain
@@ 4.1. Estimate of the $\nu$-measure of balls for general cases
In [29, Proposition 6.3], Kleinbock and Weiss proved that any $\delta$-Ahlfors regular measure with $\delta>d-1$ is absolutely $(C,\delta+1-d)$-decaying for some constant $C$ depending only on $d$ and the measure $\mu$, in the following sense: for any ball $B(\mathbf{x},r)\subset\mathbb{R}^{d}$, any hyperplane $L\subset\mathbb{R}^{d}$, and any $\varepsilon>0$,
@@
Case 3B: $Q^{-(1+1/d)}/(8d!)\leq r<Q^{-\beta\delta/(\delta+\frac{1+d}{1+\tau}-d)}$. Divide the ball $B(\mathbf{x},r)$ into smaller hypercubes of equal side length $Q^{-(1+1/d)}/(8d!)$. If any such hypercube $D$ has a nonempty intersection with $K$, then $2D\cap K$ contains a ball of radius $|D|$, where $2D$ denotes the hypercube with the same center as $D$ but twice the side length. Since any point in $K$ belongs to at most $3^{d}$ different sets of the form $2D\cap K$, a simple volume argument shows that the ball $B(\mathbf{x},r)$ can intersect at most
@@
hypercubes $D$. Furthermore, since each $D$ is convex and its $d$-dimensional Lebesgue measure is bounded by $Q^{-(1+d)}/(8d!)^{d}\leq(2Q)^{-(1+d)}/(d!)$, it follows from the simplex lemma above (see Lemma 4.1) that the rational points with denominator $1\leq q\leq 2Q$ contained in $D$ lie in some hyperplane $L\subset\mathbb{R}^{d}$. Therefore,
(4.8) $D\cap A_{Q}(cQ^{-\tau},\mathbf{0})\subset D\cap L^{(cQ^{-(1+\tau)})}.$
@@
Case 4B: $cQ^{-(1+\tau)}\leq r<Q^{-(1+1/d)}/(8d!)$. Since the $d$-dimensional Lebesgue measure of the ball $B(\mathbf{x},r)$ is less than $Q^{-(1+d)}/(8d!)^{d}$, by the same reason as (4.8),
@@
Case 5B: $0<r<cQ^{-(1+\tau)}$. In this case, we have
@@
$\frac{1-\delta^{2}-\beta\delta}{(1-\delta)^{2}}\leq\frac{1-\delta^{2}-(2-2\delta)\delta}{(1-\delta)^{2}}=1<\tau,$
@@ 6. Some Examples
Let $\Lambda^{*}=\bigcup_{k\geq 1}\Lambda^{k}$ denote the set of finite words over $\Lambda$. Let $k\geq 1$. Given a word $\omega=(i_{1},\dots,i_{k})\in\Lambda^{k}$, we define
(6.3) $\begin{split}&\phi_{\omega}:=\phi_{i_{1}}\circ\cdots\circ\phi_{i_{k}},\quad\rho_{\omega}:=\rho_{i_{1}}\cdots\rho_{i_{k}},\\ &K_{\omega}:=\phi_{\omega}(K)\quad\text{and}\quad\mu_{\omega}:=\frac{\mu|_{K_{\omega}}}{\mu(K_{\omega})}=\mu\circ\phi_{\omega}^{-1}.\end{split}$
We call $\mu_{\omega}$ a branch of $\mu$. Intuitively, $\mu_{\omega}$ represents the measure $\mu$ restricted to a “cylinder” set $K_{\omega}$ and then rescaled to a probability measure.
@@
For $t\in\mathbb{R}$ and $\mathbf{x}\in\mathbb{R}^{d}$, let
(6.4) $g(t)=\begin{pmatrix}e^{t/d}\mathbb{I}_{d}&0\\ 0&e^{-t}\end{pmatrix}\mathrm{\quad and\quad}u(\mathbf{x})=\begin{pmatrix}\mathbb{I}_{d}&\mathbf{x}\\ 0&1\end{pmatrix}$,
@@
Corollary 6.2 ([12, Corollary 2.6]).
Let $K$ be a strongly irreducible self-simlar set of an IFS $\mathcal{F}$ satisfying the OSC and let $\mu$ be the corresponding $\delta$-Ahlfors regular self-similar measure. Let $f\in B_{p,l}^{\infty}(X)$. Then, for any $\omega\in\Lambda^{*}$ and $y\in X$, we have
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\omega}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|\ll\mathrm{inj}(y)^{-1}\rho_{\omega}^{-1-\frac{d\kappa}{d+1}}\mathcal{S}_{p,l}(f)e^{-\kappa t},$
@@
Since $\mu_{\omega}=\mu\circ\phi_{\omega}^{-1}$ and $\phi_{\omega}(\mathbf{x})=\rho_{\omega}\mathbf{x}+\mathbf{b}_{\omega}$ with $\mathbf{b}_{\omega}=\phi_{\omega}(\bf 0)$, we have
$\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\omega}(\mathbf{x})$
@@
$=$ $\int_{\mathbb{R}^{d}}f\left(g\left(t+\frac{d}{d+1}\ln\rho_{\omega}\right)u(\mathbf{x})g\left(-\frac{d}{d+1}\ln\rho_{\omega}\right)u(\mathbf{b}_{\omega})y\right){\,\mathrm{d}}\mu(\mathbf{x}),$
@@
$\bigg|\int_{\mathbb{R}^{d}}f(g(t)u(\mathbf{x})y){\,\mathrm{d}}\mu_{\omega}(\mathbf{x})-\int_{X}f{\,\mathrm{d}}m_{X}\bigg|$
@@
$\mathrm{inj}(hz)\geq C_{1}\,\|\mathrm{Ad}(h)\|_{\infty}^{-1}\,\mathrm{inj}(z),$
where $\mathrm{Ad}$ denotes the adjoint action of $G$ on $\mathfrak{g}$ and $\|\cdot\|_{\infty}$ is the operator norm. Applying this with $h=g(-\frac{d}{d+1}\ln\rho_{\omega})u(\mathbf{b}_{\omega})$ yields
$\begin{split}\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\omega}\right)u(\mathbf{b}_{\omega})y\right)&\geq C_{1}\left\|\operatorname{Ad}\left(g\left(-\frac{d}{d+1}\ln\rho_{\omega}\right)\right)\right\|_{\infty}^{-1}\mathrm{inj}(u(\mathbf{b}_{\omega})y)\\ &\geq C_{1}\,\rho_{\omega}\cdot\mathrm{inj}(u(\mathbf{b}_{\omega})y).\end{split}$
Since $u(\mathbf{b}_{\omega})$ is uniformly bounded over $\omega\in\Lambda^{*}$, there exists a constant $C_{2}>0$ such that $\mathrm{inj}(u(\mathbf{b}_{\omega})y)\geq C_{2}\,\mathrm{inj}(y)$. Combining these bounds, we obtain
$\mathrm{inj}\left(g\left(-\frac{d}{d+1}\ln\rho_{\omega}\right)u(\mathbf{b}_{\omega})y\right)^{-1}\leq C_{1}^{-1}C_{2}^{-1}\rho_{\omega}^{-1}\mathrm{inj}(y)^{-1},$
@@
Lemma 6.3 ([25, Theorem 9.1 and Lemma 12.7]).